5) A pulsed solid state laser operating at wavelength λ= 1 μm and pulse duration of20 ns is used for material processing application with a focusing lens of f-number-1.8. Calculate the pulse energy required to produce an intensity 1- 2 x100 Wicm? at the focusing point. Laser physics

5th edition
Answer #1

Intensity of the laser = 2 x 1010 W/cm2 = 2 x 1014 W/m2

now, using the diffraction limit

d\frac{x}{f} = 1.22 \lambda

here, d/f = 1/1.8

so, x = spot size = 2.196 x 10-6 m

the Power deposited at the focal spot will thus be:

P= I A = 2\times 10^{14}\times \pi (2.196\times 10^{-6}/2)^2 = 757.503 W

pulse duration = 20 ns

=> Energy deposited = P x pulse duration = 757.503 x 20ns = 1.515 x 10-5 J.